# OptiHive: Ensemble Selection for LLM-Based Optimization via Statistical Modeling

### v1

@@ Introduction
where $S_{j}$ is the set of emitters within range of client $j$. This makes the problem non-linear and intractable, but rearranging the terms and taking the log on both sides allows to formulate equivalent but linear equalities
@@
which reduces the problem to an LP.
@@
The results show that across all variants, OptiHive significantly improves both optimality and feasibility rates over the baseline generator, even when component quality degrades. This demonstrates that the latent-class model can recover signal from imperfect and noisy components. In the $K$-robust and probabilistic variants, reducing the quality of the set of instances or tests has little effect on performance. This is consistent with the fact that the complexity of these variants stems from conceptual depth rather than substantial coding effort, and supports our hypothesis that tests are generally substantially easier to write correctly than the solvers themselves.
In contrast, the time-dependent variant suffers a substantial performance drop when the quality of either instances or tests is weakened. Generating a balanced set of feasible and infeasible instances is harder in this setting, and the weaker model produces an overwhelming number of infeasible instances instead of the desirable, more even split, reducing the signal available for latent inference. Moreover, this variant exposes a scenario where validity tests are themselves difficult to produce. This challenges a core assumption of our method: that testing a solution is typically easier than generating one, allowing OptiHive to differentiate good solvers through test signals. Remarkably, even in this setting, tests generated by the small model still provide a noisy but meaningful signal that allows OptiHive to outperform the baseline. This highlights OptiHive’s ability to extract value from an ensemble of entirely noisy components a crucial feature for LLM-based optimization.
Lastly, OptiHive can successfully identify optimal solvers even when most candidates are infeasible. For the time-dependent WSCP, it raise the optimality rate from 3% to 64.1%. Note that with the choice $N_{S}=50$ solvers per run for this experiment, perfect selection would achieve a 78.2% optimality rate. Conditioned on sampling at least one optimal solver in the pool of 50, we observe that OptiHive successfully selects an optimal solver 81.6% of the time in the reference setting, and still obtains a 30.9% and 25.3% success rates with degraded instances and degraded tests, respectively.
@@ Conclusion
We introduced OptiHive, a two-stage framework that (i) generates candidate solvers, instances, and tests in parallel, evaluates, and filters out unusable components, and (ii) fits a latent-class model to infer the true quality of solvers, enabling calibrated solver selection. By efficiently extracting high-performing solvers from a pool of imperfect candidates, OptiHive can either deliver high-quality solvers with minimal latency, or enhance existing solver-generation pipelines by acting as a wrapper.
Empirical results show that OptiHive substantially improves performance on challenging optimization tasks. It raises the optimality rate on simple problems to near perfection and recovers underrepresented optimal solvers on the hardest problems, increasing optimality from 5% to 92%. Our ablation study further reveals that the quality of each component type - solvers, instances, and tests - plays a significant role in overall effectiveness. In particular, high-quality instances and tests provide a crucial signal for distinguishing optimal solvers. Still, even when components are produced by smaller models, OptiHive consistently delivers improved performance over baselines.

### v2

@@ Introduction
where $S_{j}$ is the set of emitters within range of client $j$. This makes the problem non-linear and intractable, but rearranging the terms and taking the log on both sides allows us to formulate equivalent but linear inequalities
@@
which reduces the problem to an ILP.
@@
The results show that across all variants, OptiHive significantly improves both optimality and feasibility rates over the baseline generator, even when component quality degrades. This demonstrates that the latent-class model can recover signal from imperfect and noisy components. In the $K$-robust and probabilistic variants, reducing the quality of the set of instances or tests has little effect on performance. This supports our hypothesis that tests are generally substantially easier to write correctly than the solvers themselves.
In contrast, the time-dependent variant suffers a substantial performance drop when the quality of either instances or tests is weakened. Generating a balanced set of feasible and infeasible instances is harder in this setting, and the weaker model produces an overwhelming number of infeasible instances instead of a more even split. Moreover, this variant exposes a scenario where validity tests are themselves difficult to produce. This challenges a core assumption of our method: that testing a solution is typically easier than generating one, allowing OptiHive to differentiate good solvers through test signals. Remarkably, even in this setting, tests generated by the small model still provide a noisy but meaningful signal that allows OptiHive to outperform the baseline. This highlights OptiHive’s ability to extract value from an ensemble of entirely noisy components, a crucial feature for LLM-based optimization.
Lastly, OptiHive can successfully identify optimal solvers even when most candidates are infeasible. For the time-dependent WSCP, it raises the optimality rate from 3% to 64.1%. Note that with the choice $N_{S}=50$ solvers per run for this experiment, perfect selection would achieve a 78.2% optimality rate. Conditioned on sampling at least one optimal solver in the pool of 50, we observe that OptiHive successfully selects an optimal solver 81.6% of the time in the reference setting, and still obtains 30.9% and 25.3% success rates with degraded instances and degraded tests, respectively.
@@ Conclusion
We introduced OptiHive, a two-stage framework that (i) generates candidate solvers, instances, and tests in parallel, evaluates, and filters out unusable components, and (ii) fits a latent-class model to infer the true quality of solvers, enabling calibrated solver selection. By efficiently extracting high-performing solvers from imperfect candidate pools, OptiHive wraps around any solver-generation pipeline to deliver higher-quality solvers with negligible added latency.
Empirical results show that OptiHive substantially improves performance on challenging optimization tasks. It raises the optimality rate on simple problems to near perfection and recovers underrepresented optimal solvers on the hardest problems, increasing optimality from 5% to 92%. Our ablation study further reveals that the quality of each component type – solvers, instances, and tests – plays a significant role in overall effectiveness. In particular, high-quality instances and tests provide a crucial signal for distinguishing optimal solvers. Still, even when components are produced by smaller models, OptiHive consistently delivers improved performance over baselines.
@@ Acknowledgment
A  Prompt Templates
A.1 Solvers
The following prompt is used to produce solvers given a problem description, input template, and output template.
Solver Generation You are a code-generation agent expert in Python and Gurobi. [Problem Specifications]
Here is the problem description: {problem_description}
Here is the template for the problem input: {input_template} The output of the function must follow the template: {output_template} [Instructions]
Your task is to implement a function solve with a unique argument data as input and returning a solution to the problem.
Write the complete, executable, and well-indented code of the solve function, including necessary imports. Status codes are: OPTIMAL for a proven best feasible solution, INFEASIBLE when no feasible solution is found. Use a TimeLimit of 5 seconds for the optimization. Do not include example usage.
A.2 Instances
We encourage instance diversity via two mechanisms. First, we rotate diversity directives among six options:
  1. 1.
    If possible, the data should be an infeasible instance for the above problem.
  2. 2.
    The data should be a clearly feasible instance for the above problem in that it admits a simple feasible solution.
  3. 3.
    The data should result in optimal solutions to the above problem having tight constraints.
  4. 4.
    The data should be randomized.
  5. 5.
    The data should be randomized with hyperparameters that will make the instance likely feasible.
  6. 6.
    The data should be randomized with hyperparameters that will make the instance likely infeasible.
Second, we sample and provide a random seed to limit the similarity of numerical values across generated instances. To avoid lengthy LLM outputs, we ask the LLM to provide a function that returns an instance of the considered problem, rather than directly providing an instance.
We implement these guidelines through the prompt below.
Instance Generation You are a code‐generation agent expert in Python. [Problem Specifications]
Consider the following problem: {problem_description}
Here is the template for the problem input: {input_template} [Instructions]
Your task is to implement a function generate_input with no argument and returning a input following the input template.
{diversity_directives} Write the complete, executable and well indented code of the generate_input function, including necessary imports. Use the following seed: {seed}
A.3 Tests
Test Generation You are a code‐generation agent expert in Python. [Problem Specifications]
Here is the problem description: {problem_description}
Here is the input template: {input_template} Here is the solution template: {output_template} [Instructions]
For every concrete instance data that follows the input template, there is a corresponding solution object that follows the solution template.
Your task is to implement a function test(data, solution) -> bool that returns True if and only if all of the following hold: 1. The solution is feasible (it satisfies every problem constraint). 2. The reported objective value matches the cost you compute (within a small numerical tolerance). 3. All solution fields are internally coherent. Write the complete, executable, and well‑indented Python code implementing the test function, including necessary imports. Do not include example usage.
B  Expectation Maximization Model Specifications
Observed variables:
$r_{s,i}$ $=\left\{\begin{tabular}[]{cl}1&\text{if solver $s$ reports a solution}\\ 0&\text{otherwise.}\end{tabular}\right.,\quad\forall(s,i)\in S\times I$
$r_{s,i,t}$ $=\left\{\begin{tabular}[]{cl}1&\text{solution $(s,i)$ passes test $t$}\\ 0&\text{otherwise.}\end{tabular}\right.,\quad\forall(s,i,t)\in S\times I\times T\text{ s.t. }r_{s,i}=1$
For all $(s,i)\in S\times I$ such that $r_{s,i}=1$, we define $C_{s,i}=\sum_{t\in T}r_{s,i,t}$.
Latent variables:
$f_{i}$ $=\left\{\begin{tabular}[]{cl}1&\text{if instance $i$ admits a feasible solution}\\ 0&\text{otherwise.}\end{tabular}\right.,\quad\forall i\in I$
$f_{s,i}$ $=\left\{\begin{tabular}[]{cl}1&\text{solution $(s,i)$ is feasible}\\ 0&\text{otherwise.}\end{tabular}\right.,\quad\forall(s,i)\in S\times I\text{ s.t. }r_{s,i}=1$
Parameters:
$\theta=\left(\lambda,(\alpha_{s})_{s\in S},(\beta_{s})_{s\in S},(\gamma_{s})_{s\in S},a_{0},b_{0},a_{1},b_{1}\right)$
where
$\lambda$ $=\mathbb{P}\left(f_{i}=1\right)$ (feasiblity rate of instances)
$\alpha_{s}$ $=\mathbb{P}\left(r_{s,i}=1\mid f_{i}=0\right)$ (type I error of solver $s$)
$\beta_{s}$ $=\mathbb{P}\left(r_{s,i}=0\mid f_{i}=1\right)$ (type II error of solver $s$)
$\gamma_{s}$ $=\mathbb{P}\left(f_{s,i}=1\mid f_{i}=1,r_{s,i}=1\right)\quad$ (feasibility rate of solutions reported by $s$ on feasible instances)
$a_{0},b_{0}$ Parameters of the Beta–Binomial for infeasible solutions
$a_{1},b_{1}$ Parameters of the Beta–Binomial for feasible solutions
B.1 E-step
We first compute the intermediate quantities involved in the conditional expectations of $f_{i}$ and $f_{s,i}$. Let R denote all observed variable and $\textbf{R}_{s}=\{r_{s,i}:\;i\in I\}\cup\left\{r_{s,i,t}:\;(i,t)\in I\times T,r_{s,i}=1\right\}$ denote the observed variables related to solver $s$. Let $\nu(k\mid n,a,b)=\binom{n}{k}\frac{B(k+a,n-k+b)}{B(a,b)}$ be the probability mass function of the Beta-Binomial distribution, where $B$ is the beta function. We define for all $(s,i)\in S\times I$:
$A_{s,i}^{(0)}$ $\triangleq\nu(C_{s,i}\mid T,a_{0},b_{0})$
$A_{s,i}^{(1)}$ $\triangleq\nu(C_{s,i}\mid T,a_{1},b_{1})$
and, for all $i\in I$:
$B_{i}^{(1)}$ $\triangleq\mathbb{P}\left(\textbf{R}\mid f_{i}=1,\mathbf{\theta}\right)$
$=\prod_{s\in S}\left[(1-\beta_{s})\left(\gamma_{s}\mathbb{P}\left(\textbf{R}_{s}\mid r_{s,i}=1,f_{s,i}=1,\mathbf{\theta}\right)+(1-\gamma_{s})\mathbb{P}\left(\textbf{R}_{s}\mid r_{s,i}=1,f_{s,i}=0,\mathbf{\theta}\right)\right)\right]^{r_{s,i}}\beta_{s}^{(1-r_{s,i})}$
$=\prod_{s\in S}\left[(1-\beta_{s})\left(\gamma_{s}A_{s,i}^{(1)}+(1-\gamma_{s})A_{s,i}^{(0)}\right)\right]^{r_{s,i}}\beta_{s}^{(1-r_{s,i})}$
$B_{i}^{(0)}$ $\triangleq\mathbb{P}\left(\textbf{R}\mid f_{i}=0,\mathbf{\theta}\right)$
$=\prod_{s\in S}\left[\alpha_{s}\mathbb{P}\left(\textbf{R}_{s}\mid f_{i}=0,\mathbf{\theta}\right)\right]^{r_{s,i}}(1-\alpha_{s})^{(1-r_{s,i})}$
$=\prod_{s\in S}\left[\alpha_{s}A_{s,i}^{(0)}\right]^{r_{s,i}}(1-\alpha_{s})^{(1-r_{s,i})}$
We then proceed to compute the conditional expectation of $f_{i}$ and $f_{s,i}$ for a given $\theta$. For all $i\in I$:
$\mathbb{E}\left[f_{i}\mid\textbf{R},\mathbf{\theta}\right]$ $=\mathbb{P}\left(f_{i}=1\mid\textbf{R},\mathbf{\theta}\right)$
$=\frac{\mathbb{P}\left(\textbf{R}\mid f_{i}=1,\mathbf{\theta}\right)\mathbb{P}\left(f_{i}=1\mid\mathbf{\theta}\right)}{\mathbb{P}\left(\textbf{R}\mid\mathbf{\theta}\right)}$
$=\frac{\mathbb{P}\left(\textbf{R}\mid f_{i}=1,\mathbf{\theta}\right)\mathbb{P}\left(f_{i}=1\mid\mathbf{\theta}\right)}{\mathbb{P}\left(\textbf{R}\mid f_{i}=1,\mathbf{\theta}\right)\mathbb{P}\left(f_{i}=1\right)+\mathbb{P}\left(\textbf{R}\mid f_{i}=0,\mathbf{\theta}\right)\mathbb{P}\left(f_{i}=0\right)}$
$=\frac{\lambda B_{i}^{(1)}}{\lambda B_{i}^{(1)}+(1-\lambda)B_{i}^{(0)}}$
For all $(s,i)\in S\times I$ such that $r_{s,i}=1$:
$\mathbb{P}\left(f_{s,i}=1\mid f_{i}=1,\textbf{R},\mathbf{\theta}\right)$
$\quad=\frac{\mathbb{P}\left(\textbf{R}\mid f_{s,i}=1,f_{i}=1,\mathbf{\theta}\right)\mathbb{P}\left(f_{s,i}=1\mid f_{i}=1,\mathbf{\theta}\right)}{\mathbb{P}\left(\textbf{R}\mid f_{i}=1,\mathbf{\theta}\right)}$
$\quad=\frac{\mathbb{P}\left(\textbf{R}\mid f_{s,i}=1,f_{i}=1,\mathbf{\theta}\right)\mathbb{P}\left(f_{s,i}=1\mid f_{i}=1,\mathbf{\theta}\right)}{\mathbb{P}\left(\textbf{R}\mid f_{s,i}=1,f_{i}=1,\mathbf{\theta}\right)\mathbb{P}\left(f_{s,i}=1\mid f_{i}=1\right)+\mathbb{P}\left(\textbf{R}\mid f_{s,i}=0,f_{i}=1,\mathbf{\theta}\right)\mathbb{P}\left(f_{i}=0\mid f_{i}=1\right)}$
$\quad=\frac{\gamma_{s}A_{s,i}^{(1)}}{\gamma_{s}A_{s,i}^{(1)}+(1-\gamma_{s})A_{s,i}^{(0)}}$
Since $\mathbb{P}\left(f_{s,i}=1\mid f_{i}=0\right)=0$, we obtain:
$\mathbb{E}\left[f_{s,i}\mid\textbf{R},\mathbf{\theta}\right]$ $=\mathbb{P}\left(f_{s,i}=1\mid f_{i}=1,\textbf{R},\mathbf{\theta}\right)\mathbb{P}\left(f_{i}=1\mid\textbf{R},\mathbf{\theta}\right)+\mathbb{P}\left(f_{s,i}=1\mid f_{i}=0,\textbf{R},\mathbf{\theta}\right)\mathbb{P}\left(f_{i}=0\mid\textbf{R},\mathbf{\theta}\right)$
$=\frac{\gamma_{s}A_{s,i}^{(1)}}{\gamma_{s}A_{s,i}^{(1)}+(1-\gamma_{s})A_{s,i}^{(0)}}\frac{\lambda B_{i}^{(1)}}{\lambda B_{i}^{(1)}+(1-\lambda)B_{i}^{(0)}}$
B.2 M-step
The log-likelihood is given by:
$\ln$ $\;\mathbb{P}\left(\textbf{R},\textbf{F}\mid\mathbf{\theta}\right)$
$=\sum_{i\in I}\left[\ln\mathbb{P}\left(f_{i}\right)+\sum_{s\in S}\ln\mathbb{P}\left(r_{s,i}\mid f_{i}\right)+\sum_{\begin{subarray}{c}s\in S\\ r_{s,i}=1\end{subarray}}\left[\ln\mathbb{P}\left(f_{s,i}\mid f_{i}\right)+\sum_{t\in T}\ln\mathbb{P}\left(r_{s,i,t}\mid f_{s,i}\right)\right]\right]$
$=\sum_{i\in I}\left[f_{i}\ln(\lambda)+(1-f_{i})\ln(1-\lambda)\right]$
$\quad+\sum_{i\in I}\sum_{s\in S}\left[r_{s,i}\left(f_{i}\ln(1-\beta_{s})+(1-f_{i})\ln(\alpha_{s})\right)+(1-r_{s,i})\left(f_{i}\ln(\beta_{s})+(1-f_{i})\ln(1-\alpha_{s})\right)\right]$
$\quad+\sum_{i\in I}\sum_{\begin{subarray}{c}s\in S\\ r_{s,i}=1\end{subarray}}f_{i}\left[f_{s,i}\ln(\gamma_{s})+(1-f_{s,i})\ln(1-\gamma_{s})\right]$
$\quad+\sum_{i\in I}\sum_{\begin{subarray}{c}s\in S\\ r_{s,i}=1\end{subarray}}\left[f_{s,i}\ln\nu(C_{s,i}\mid T,a_{1},b_{1})+(1-f_{s,i})\ln\nu(C_{s,i}\mid T,a_{0},b_{0})\right]$
Let $\hat{f}_{i}=\mathbb{E}\left[f_{i}\mid\textbf{R},\theta\right]$ and $\hat{f}_{s,i}=\mathbb{E}\left[f_{s,i}\mid\textbf{R},\theta\right]=\mathbb{E}\left[f_{i}f_{s,i}\mid\textbf{R},\theta\right]$. Taking the conditional expectation and differentiating with respect to $\lambda$, we have:
$\frac{\partial}{\partial\lambda}\mathbb{E}\left[\ln\mathbb{P}\left(\textbf{R}\mid\mathbf{\theta}\right)\right]=\sum_{i\in I}\left(\frac{\hat{f}_{i}}{\lambda}-\frac{1-\hat{f}_{i}}{1-\lambda}\right)$ (4)
Equating (4) to 0, we obtain: $\lambda=\frac{\sum_{i\in I}\hat{f}_{i}}{I|}$. With analogous reasoning, we have:
$\alpha_{s}$ $=\frac{\sum_{i\in I}(1-\hat{f}_{i})r_{s,i}}{\sum_{i\in I}(1-\hat{f}_{i})}$
$\beta_{s}$ $=\frac{\sum_{i\in I}\hat{f}_{i}(1-r_{s,i})}{\sum_{i\in I}\hat{f}_{i}}$
$\gamma_{s}$ $=\frac{\sum_{i\in I}r_{s,i}\hat{f}_{s,i}}{\sum_{i\in I}r_{s,i}\hat{f}_{i}}$
We use the method of moments to update the parameters $a_{0}$, $b_{0}$, $a_{1}$, and $b_{1}$ using a Beta prior with $(\bar{\alpha},\bar{\beta})=(20,1)$ for the true positive rate. Define:
$p_{0}$ $=\frac{\sum_{s,i}\left(1-\hat{f}_{s,i}\right)\frac{C_{s,i}}{T}}{\sum_{s,i}\left(1-\hat{f}_{s,i}\right)}$
$p_{1}$ $=\frac{\sum_{s,i}\hat{f}_{s,i}\frac{C_{s,i}}{T}+\bar{\alpha}-1}{\sum_{s,i}\hat{f}_{s,i}+\bar{\alpha}+\bar{\beta}-2}$
and
$\mu_{0}$ $=\frac{\sum_{s,i}\left(1-\hat{f}_{s,i}\right)C_{s,i}}{T\cdot\sum_{s,i}\left(1-\hat{f}_{s,i}\right)}$
$\sigma^{2}_{0}$ $=\frac{\sum_{s,i}\left(1-\hat{f}_{s,i}\right)\left(C_{s,i}/T-\mu_{0}\right)^{2}}{\sum_{s,i}\left(1-\hat{f}_{s,i}\right)}$
$\mu_{1}$ $=\frac{\sum_{s,i}\hat{f}_{s,i}C_{s,i}}{T\cdot\sum_{s,i}\hat{f}_{s,i}}$
$\sigma^{2}_{1}$ $=\frac{\sum_{s,i}\hat{f}_{s,i}\left(C_{s,i}/T-\mu_{1}\right)^{2}}{\sum_{s,i}\hat{f}_{s,i}}$
Then for $k\in\{0,1\}$:
$\rho_{k}=\frac{\frac{T\sigma_{k}^{2}}{\mu_{k}(1-\mu_{k})}-1}{T-1}$
Finally, retrieve $a_{k}$ and $b_{k}$ from $p_{k}$ and $\rho_{k}$ for $k\in\{0,1\}$ as:
$a_{k}$ $=p_{k}\left(\frac{1}{\rho_{k}}-1\right)$
$b_{k}$ $=(1-p_{k})\left(\frac{1}{\rho_{k}}-1\right)$